5. Let w = i.$= (4.2+3.4.2-2w). 1 = (1.3w. 20.-w). Find the point wise multiplication ext.
Answer #1

5 W Given that S =(4, 2+36, 4,2-2W) +=(1,367, 269-w) : Where wai wawa ixè = ; ²=-1. * we=a1 The point wise multiplication of